A geological study indicates that an exploratory oil well should strike oil with probability
0.2.
(a) What is the probability that the rst strike comes on the third well drilled?
(b) What assumptions did you make to obtain the answer in (a).
1
Expert's answer
2020-04-16T18:00:58-0400

probability of strike oil in oil well "(P)=0.2"

probability of not strike oil in oil well "(1-P)=0.8"


(a)The probability that the first strike comes on the third well drilled"(P_3),"

"P_3=(1-P)(1-P)P=0.8*0.8*0.2=\bold{0.128}"

(b)

  • Drilling each well is independent from others.
  • Assume there are only two possibility when drilled the oil well such that oil strike or not strike.